A system of equations is given below, 𝑑π‘₯ + 2𝑦 + 3𝑧 = π‘Ž 2π‘₯ + 3𝑦 βˆ’ 𝑑𝑧 = 𝑏 3π‘₯ + 5𝑦 + (𝑑 + 1)𝑧 = 𝑐 Where 𝑑 is an integer and π‘Ž, 𝑏, 𝑐 are real constants. The system does not have a unique solution, but it is consistent. Show that π‘Ž + 𝑏 = 𝑐


1
Expert's answer
2021-09-30T00:46:46-0400

Solution:

Given,

"𝑡𝑥 + 2𝑦 + 3𝑧 = 𝑎 
\\2𝑥 + 3𝑦 − 𝑡𝑧 = 𝑏 
\\3𝑥 + 5𝑦 + (𝑡 + 1)𝑧 = 𝑐"

The system does not have a unique solution, but it is consistent.

It means the system has infinite number of solutions.

Augmented matrix is "\begin{bmatrix}t&2&3&|a\\ 2&3&-t&|b\\ 3&5&t+1&|c\end{bmatrix}"

We need to make all elements of row 3 zero as the system has infinite number of solutions.

Reduce matrix to row echelon form.

"=\begin{bmatrix}t&2&3&a\\ 0&\frac{5t-6}{t}&\frac{t^2+t-9}{t}&\frac{tc-3a}{t}\\ 0&0&\frac{-8t^2+7t+1}{5t-6}&\frac{5tb-3tc-a-6b+4c}{5t-6}\end{bmatrix}"

Now, we must have "\dfrac{-8t^2+7t+1}{5t-6}=0" and "\dfrac{5tb-3tc-a-6b+4c}{5t-6}=0" ...(i)

From "\dfrac{-8t^2+7t+1}{5t-6}=0"

"\Rightarrow -8t^2+7t+1=0
\\ \Rightarrow t=1, t=-\dfrac18"

Given that t is an integer, so t must be equal to 1, not "-\dfrac18" .

Put t=1 in (i)

"\dfrac{5b-3c-a-6b+4c}{5-6}=0
\\\Rightarrow -b-a+c=0
\\\Rightarrow a+b=c"

Hence, proved.